# Criticality-Guided Efficient Pruning in Spiking Neural Networks Inspired by Critical Brain Hypothesis

### v2

@@ 4.3.1 Unstructured Pruning
To better evaluate the effectiveness of the criticality-based regeneration mechanism, we choose the simple global magnitude pruning criterion $|w|$ to achieve connection level sparsity in deep SNNs, where $w$ is the connection weight. Following the gradual pruning scheme (details in Appendix A) in (Zhu and Gupta, 2017), we iteratively prune the network until reaching the final target sparsity and perform the regeneration based on neuron criticality after each pruning iteration. Specifically, we collect the values of $g^{\prime}(u)$ from the last training iteration before pruning and calculate the average criticality score. Then, we perform the global regeneration after pruning.
@@ 4.3.2 Structured Pruning
We employ the scaling factor of the batch normalization layer to obtain channel-level sparse models following (Liu et al., 2017) (details in Appendix A). Specifically, we leverage the parameter $\gamma$ in batch normalization layers as the pruning criterion and incorporate L1 sparsity regularization during training. After training, we prune channels according to the magnitude of $\gamma$. Then, we calculate the criticality score for each channel by collecting the average values of the surrogate function derivative from the entire training dataset. Finally, we execute regeneration with criticality scores and fine-tune the pruned network.
@@ 5.1 Experimental Setup
We evaluate our method on a shallow SNN(6-conv-2-fc) and two representative deep SNNs VGG-16 (Simonyan and Zisserman, 2014) and ResNet-19 (He et al., 2016) for three datasets CIFAR10, CIFAR100 (Krizhevsky et al., 2009) and Tiny-ImageNet (Hansen, 2015). We use batch size 128, SGD optimizer with momentum 0.9 and learning rate 0.3. We set 200 epochs for iterative pruning for CIFAR10 and CIFAR100 (100 for tiny-ImageNet) and the timestep to 5 for all experiments. Following previous works, we use batch normalization (BN) and take the first convolutional layer as the encoder to transform the input images into spikes. Our implementation is based on PyTorch. The experiments are executed on an RTX 3090 GPU. More implementation details are contained in Appendix B.
@@ 5.2 Accuracy and Efficiency Evaluations
Pruning Method Dataset Arch. Base Acc. (%) Cost Pruned Acc. (%) Sparsity (%)
Grad R (Chen et al., 2021) CIFAR10 6Conv, 2Fc 92.84 54.61 91.47 97.65
@@
RCMO-SNN (Chen et al., 2023) CIFAR10 6Conv, 2Fc 92.88 54.61 92.75 97.71
@@
Ours CIFAR10 6Conv, 2Fc 92.79 3.33 92.77$\pm$0.21 97.71
@@
NDSNN (Huang et al., 2023) CIFAR100 VGG16 69.86 5.00 68.07 90.00
@@
Grad R (Chen et al., 2021) CIFAR100 ResNet19 71.34 34.13 67.47 94.92
@@
IMP (Kim et al., 2022) CIFAR100 ResNet19 71.34 37.67 71.38 89.91
56.33 70.54 95.69
70.33 67.35 98.13
RCMO-SNN (Chen et al., 2023) CIFAR100 ResNet19 74.71 51.67 72.67 95.19
65.67 70.80 97.31
Ours CIFAR100 ResNet19 74.31 3.33 73.65$\pm$0.23 90.00
@@
RigL (Evci et al., 2020) Tiny-ImageNet ResNet19 50.32 7.50 49.49 90.00
@@
NDSNN (Huang et al., 2023) Tiny-ImageNet ResNet19 50.32 7.50 49.25 90.00
@@
IMP 1  (Kim et al., 2023) Tiny-ImageNet ResNet19 58.00 2.50 54.94 90.00
@@
Ours Tiny-ImageNet ResNet19 59.85 2.50 58.00$\pm$0.16 90.00
@@
Pruning Method Pruned Acc. (%) FLOPs normal-$\downarrow$ (%)
@@
 (Liu et al., 2017) 67.92 43.76
66.24 58.10
 (Wang et al., 2021) 68.05 43.73
67.13 58.01
Ours 68.90 43.76
67.64 58.02
@@
 (Liu et al., 2017) 71.08 49.44
68.34 69.49
 (Wang et al., 2021) 70.37 49.40
69.20 69.42
Ours 71.28 49.46
70.01 69.47
Method Sparsity
90.00% 95.69% 98.13%
@@
5.2.1 Unstructured Pruning
We evaluate our method in Table 2 and compare it with state-of-the-art (SOTA) pruning methods for SNNs, including Grad R (Chen et al., 2021), STDS (Chen et al., 2022), RigL (Evci et al., 2020), IMP (Kim et al., 2022), NDSNN (Huang et al., 2023), and RCMO-SNN (Chen et al., 2023). We report accuracy and pruning costs at different sparsity levels. The pruning cost is estimated as $\frac{T\times N_{p}}{N_{d}}$ where $T$ is the number of time steps, $N_{p}$ is the total number of epochs during pruning, and $N_{d}$ is the number of epochs for training dense models in this work (same for all methods). The costs of different methods are normalized with $N_{d}$. On CIFAR10 with the shallow 6conv-2fc SNN, our method outperforms previous approaches at all sparsity levels with lower pruning costs. On the more complex datasets CIFAR100 and Tiny-ImageNet, our method exhibits even more significant improvements. As shown in Table 2, On CIFAR100, we achieve higher accuracy compared to IMP and RCMO-SNN with a pruning cost reduction ranging from 91.15% to 95.26% on VGG16 and ResNet19. For Tiny-ImageNet, our method demonstrates from 3.06% to 5.21% performance advantages over IMP with the same pruning costs.
5.2.2 Structured Pruning
In Table 3, we evaluate the performance of our method on CIFAR100 for structured pruning. As pioneers in implementing structured pruning on SNNs, we compare our method with the SOTA works of ANN pruning NSlim (Liu et al., 2017) and GReg (Wang et al., 2021) and report the pruned test accuracy and the float point operations reduction (FLOPs $\downarrow$) in inference. Additionally, we take NSlim as our baseline because our method develops based on it. Our method consistently outperforms GReg (Wang et al., 2021), which is more complex and computationally expensive, and shows significant improvements compared to the baseline, all while maintaining the equivalent level of FLOPs.
5.3 Effect of Neuron Criticality
Figure 2: Comparison of regeneration methods based on criticality (ours), gradient, random, and GraNet Liu et al. (2021). The two dashed lines represent the performance of the dense model and the baseline, respectively.
@@ 5.3 Effect of Neuron Criticality
We compare the baseline and pruning based on neuron criticality on CIFAR100. The results of the ablation study are shown in Table 4, where the proposed method outperforms the baseline at all sparsity levels. Furthermore, to evaluate the contribution of the criticality-based regeneration mechanism, we perform a performance analysis in Figure 2 using various regeneration metrics, including criticality, Grad Evci et al. (2020), Random, and GraNet (Liu et al., 2021). ”Criticality” refers to the method based on criticality, ”Grad” employs parameter gradients at the current iteration, ”Random” adopts random ranking, and ”GraNet” represents the GraNet gradient regeneration strategy. We observe that criticality-based regeneration consistently outperforms Grad and Random regeneration, highlighting our method’s superiority.
5.4 Effect of Regeneration Ratio $r$
To evaluate the robustness of the criticality-based regeneration, we conduct performance analysis in Figure 2 using different metrics with different regeneration ratios $r$. The two dashed lines represent the dense model and baseline performance, respectively. We observe that the criticality-based method exhibits stable performance within the ratio range from 0.1 to 0.6 and higher accuracy than the baseline and other methods, indicative of the robustness over the wide range of $r$. GraNet exhibits different performance dynamics compared to other methods. We attribute this to the cosine decay of the regeneration ratio, gradually decreasing the impact of GMP as training progresses.
@@ 5.4 Effect of Regeneration Ratio $r$
5.5 Effect of Surrogate Function
To evaluate the effectiveness and robustness of the criticality mechanism for different surrogate functions, we report the pruning performance of different surrogate functions on VGG16 for CIFAR100. As shown in Figure 2(a), we compare three surrogate functions, tanh, sigmoid, and arc tangent, for the dense model, baseline, and proposed method. The proposed method achieves higher performance compared to the baseline across different surrogate functions. The formulas of surrogate functions are contained in Appendix B.
6 Discussion
In order to investigate the underly impact of neuron criticality during pruning and fine-tuning, we present the experimental results and statistical findings for the pruned model in this section for more in-depth analysis.
6.1 Importance Transition
We compare the importance transition in non-overlapping surviving structures between the models obtained by our method and the baseline. To better distinguish non-overlapping structures, we focus on the results of structured pruning. In Figure 2(b), we present the means of $\gamma$ (normalized) of non-overlapping channels in the models through the fine-tuning process with L1 sparsity regularization for two methods. The results show that the channels regenerated by our method exhibit significantly higher importance than those without regeneration. Considering that the regenerated channels initially corresponded to lower $\gamma$ values and higher criticality, this result suggests the following role of our method: identifying structures with latent potential, enabling the high-criticality structures to occupy key positions and efficiently leading to higher performance gains.
@@ 6.2 Feature Extraction
The critical brain theory suggests that the critical state of the brain promotes feature extraction and transmission. We compare the difference in feature of models pruned by our method and the baseline. Figure 3(a) shows the mean of intra-cluster variance (Kiang, 2001) of the features in CIFAR100 for unstructured and structured pruning, respectively. The intra-cluster variance measures the compactness of intra-cluster sample representations (Oti et al., 2020). We extract feature maps before the fully connected classifier and normalize them to eliminate the influence of absolute values. According to Figure 3(a), our model learns more consistent feature representations in almost all classes (we further report it in Appendix C) and achieves a lower mean of variances, indicating enhanced compactness in class features. This improvement is consistent for both unstructured and structured pruning.
Furthermore, Figure 3(b) illustrates the mean of cosine similarity of the class feature map between the training and test datasets. Our model exhibits higher similarity in feature extraction across almost all classes (we further report it in Appendix C), leading to a more uniform representation of features between the train and test samples and reducing noise. Figure 3(c) and 3(d) show the accuracies of pruned models for the training and test set after unstructured and structured pruning. Our model exhibits higher test accuracy after pruning but lower train accuracy than the baseline. We also report the training and test loss of pruned models by our method and the baseline in Appendix C, and the results demonstrate a similar pattern to Figure 3(c) and 3(d). Combining the above observation, the baseline models suffer from overfitting during fine-tuning, and our method reduces this because the critical model learns consistent feature representation and reduces noise between training and test samples.

### v3

@@ 4.3.2 Unstructured Pruning
So far, we have known how to calculate the criticality scores of neurons. we perform unstructured (connection) pruning on SNNs as depicted in Fig. 1(a). Our approach employs iterative magnitude pruning and incorporates a regeneration mechanism based on neuron criticality. We choose the simple global magnitude pruning criterion $|w|$ to achieve connection level sparsity in deep SNNs. Specifically, we have:
$s(w)=|w|,$ (14)
where $w$ is the connection weight and $s(w)$ is the importance score of the connection.
Algorithm 1 The pseudo code of unstructured pruning
1:target sparsity $s_{f}$, pruning interval $\Delta t$, total period $T_{f}$ and regeneration ratio r.
2:Model weights $W$.
3:Let $n\leftarrow 0$.
4:for each training step $t$ do
5:     TrainOneStep()
6:     if $t<T_{f}$ and $(t\bmod\Delta t)==0$ then
7:         $n=n+1$
8:         $s_{t}\leftarrow$ CurrentSparsity($s_{f},n,T_{f},\Delta t$) $\triangleright$ Eq. 15
9:         $s^{\prime}_{t}\leftarrow$ ExtendSparsity($s_{t},r$) $\triangleright$ Eq. 11
10:         $W=$ Pruning($|W|,s^{\prime}_{t}$) $\triangleright$ Eq. 14
[…]
12:         $W=$ Regeneration($C(W),s^{\prime}_{t}-s_{t}$) $\triangleright$ Eq. 12
13:     end if
14:end for
To best maintain the criticality of the pruned model, we employ a gradual pruning scheme and iteratively prune the network until reaching the final target sparsity. The gradual pruning scheme is
$s_{t}=s_{f}-s_{f}(1-\frac{n\Delta t}{T_{f}})^{3},$ (15)
where $s_{f}$ represents the final target sparsity. $n$ is the current pruning iteration. $T_{f}$ is the total period of pruning. $\Delta t$ is a constant and represents the interval of pruning iterations.
We present the overall pruning strategy with regeneration in Algorithm 1. The regeneration based on neuron criticality is included in each pruning iteration. Specifically, we collect the values of $g^{\prime}(u)$ from the last training iteration before pruning and calculate the average criticality score. Then, we perform the global regeneration after pruning.
@@ 4.3.3 Structured Pruning
To evaluate the effectiveness of applying criticality in once pruning, we conduct one-shot structured pruning as depicted in Fig. 1(b). The overall pruning strategy is presented in Algorithm 2. Specifically, we leverage the parameter $\gamma$ in batch normalization layers as the pruning criterion (Liu et al., 2017) to achieve structured channel-level sparsity, and incorporate L1 sparsity regularization during training:
$L=\sum_{(x,y)}{l(f(x,W),y)}+\lambda\sum_{\gamma\in\Gamma}{|\gamma|},$ (16)
where $(x,y)$ is the input and target pair. $W$ is network weights. The first sum term denotes train loss in the backward process. The second sum term is the L1 sparsity regularization. After training, we prune channels according to the magnitude of $\gamma$. Then, we calculate the criticality score for each channel by collecting the average values of the surrogate function derivative from the entire training dataset. Finally, we execute regeneration with criticality scores and fine-tune the pruned network.
Algorithm 2 The pseudo code of structured pruning
1:target sparsity $s_{t}$, and regeneration ratio r.
2:Model channels $chs$ and scaling factor set $\Gamma$. Train process:
3:Train model with L1 sparsity regularization. $\triangleright$ Ep.16 Pruning and Regeneration:
4:$s^{\prime}_{t}\leftarrow$ ExtendSparsity($s_{t},r$) $\triangleright$ Eq. 11
5:$chs=$ Pruning($chs,s^{\prime}_{t},|\Gamma|$) $\triangleright$ pruning by $\gamma$
[…]
7:$chs=$ Regeneration($C(chs),s^{\prime}_{t}-s_{t}$) $\triangleright$ Eq. 12 Fine-tuning process:
8:Fine-tuning model without L1 sparsity regularization.
@@ 5.1 Experimental Setup
We evaluate our method on a shallow SNN(6-conv-2-fc) and two deep SNNs VGG16 (Simonyan and Zisserman, 2014) and ResNet19 (He et al., 2016) for three image classification datasets CIFAR10, CIFAR100 (Krizhevsky et al., 2009) and Tiny-ImageNet (Hansen, 2015). All experiments are executed on an RTX 3090 GPU and the implementation is based on PyTorch. The networks are trained using SGD optimizer with momentum 0.9, learning rate 0.3 and batch size 128. We set 200 epochs for unstructured pruning for CIFAR10 and CIFAR100 (100 for tiny-ImageNet) and 160 epochs for structured pruning. To be consistent with previous works, the time step is set to 5 for all experiments, and batch normalization (BN) is used with the first convolutional layer serving as the encoder to transform input images into spikes. A step lrscheduler is used for structured pruning, while a cosine lrscheduler is applied for unstructured pruning. The detail settings are illustrated in Table 2.
Table 2: The detailed settings of hyperparameters for all experiments.
Parameter Description Value
Global Setting
$T$ Timestep 5
$\tau$ Membrane constant $4/3$
$V_{\text{threshold}}$ Threshold voltage 1.0
$V_{\text{reset}}$ Reset voltage 0.0
For Unstructured Pruning
$wd$ Weight decay 1e-4
$\Delta t$ Pruning interval 2000
$s_{f}$ Final sparsity [0.90, 0.95, 0.98]
$r$ Regeneration ratio [0.2, 0.1, 0.01]
For Structured Pruning
$N_{1},N_{2}$ Drop (10x) epochs [80, 120]
$\gamma$ Scaling factor of L1 1e-4
$percent^{1}$ Target sparsity [0.512, 0.658]
$r$ Regeneration ratio [0.3, 0.1]
  • 1
    Choosing special values to achieve the same flops reduction as SOTA for fair comparison.
@@ 5.2 Accuracy and Efficiency Evaluations
Pruning Method Dataset Arch. Base Acc. (%) #Epoch Pruned Acc. (%) Sparsity (%)
Grad R (Chen et al., 2021) CIFAR10 6Conv, 2Fc 92.84 2048 91.47 97.65
@@
RCMO-SNN (Chen et al., 2023) CIFAR10 6Conv, 2Fc 92.88 2048 92.75 97.71
@@
Ours CIFAR10 6Conv, 2Fc 92.79 200 92.77$\pm$0.21 97.71
@@
IMP (Kim et al., 2022) CIFAR100 VGG16 69.86 2260 68.90 89.91
3380 68.00 95.69
4220 66.02 98.13
NDSNN (Huang et al., 2023) CIFAR100 VGG16 69.86 300 68.07 90.00
@@
Grad R (Chen et al., 2021) CIFAR100 ResNet19 71.34 2048 67.47 94.92
@@
IMP  (Kim et al., 2022) CIFAR100 ResNet19 71.34 2860 71.38 89.91
3380 70.54 95.69
4220 67.35 98.13
RCMO-SNN (Chen et al., 2023) CIFAR100 ResNet19 74.71 3100 72.67 95.19
3940 70.80 97.31
NDSNN (Huang et al., 2023) CIFAR100 ResNet19 71.94 300 70.08 90.00
68.95 95.00
65.48 98.00
Ours CIFAR100 ResNet19 74.31 200 73.65$\pm$0.23 90.00
@@
RigL (Evci et al., 2020) Tiny-ImageNet ResNet19 50.32 300 49.49 90.00
@@
NDSNN (Huang et al., 2023) Tiny-ImageNet ResNet19 50.32 300 49.25 90.00
@@
IMP 1  (Kim et al., 2023) Tiny-ImageNet ResNet19 58.00 100 54.94 90.00
@@
Ours Tiny-ImageNet ResNet19 59.85 100 58.00$\pm$0.16 90.00
@@
Pruning Method Pruned Acc. (%) FLOPs $\downarrow$ (%)
@@
NSlim (Liu et al., 2017) 71.93 43.57
70.41 57.47
Greg (Wang et al., 2021) 72.05 43.65
70.13 57.21
Ours 72.65 43.63
71.47 57.70
@@
NSlim (Liu et al., 2017) 72.98 48.05
70.13 66.92
Greg (Wang et al., 2021) 73.45 47.10
70.88 67.69
Ours 74.63 47.26
72.51 67.85
5.2.1 Unstructured Pruning
We evaluate our method in Table 3 and compare it with state-of-the-art (SOTA) pruning methods for SNNs, including Grad R (Chen et al., 2021), STDS (Chen et al., 2022), RigL (Evci et al., 2020), IMP (Kim et al., 2022), NDSNN (Huang et al., 2023), and RCMO-SNN (Chen et al., 2023). We report accuracy and pruning costs (#Epoch) at different sparsity levels. On CIFAR10 with the shallow 6conv-2fc SNN, our method outperforms previous approaches at all sparsity levels with lower pruning costs. On the more complex datasets CIFAR100 and Tiny-ImageNet, our method exhibits even more significant improvements. As shown in Table 3, On CIFAR100, we achieve higher accuracy compared to IMP and RCMO-SNN with a pruning cost reduction ranging from 91.15% to 95.26% on VGG16 and ResNet19. For Tiny-ImageNet, our method demonstrates from 3.06% to 5.21% performance advantages over IMP with the same pruning costs. Our method also achieves smaller accuracy losses compared to nearly all SOTA methods. While IMP reports a smaller accuracy loss on ResNet19 for CIFAR100, it is based on a lower base accuracy. Chen et al. (2023) reproduced IMP, achieving a base accuracy that is 3.37% higher than the reported results in (Kim et al., 2023). Our reproduction confirms this, showing that IMP incurs greater accuracy losses.
5.2.2 Structured Pruning
In Table 4, we evaluate the performance of our method on CIFAR100 for structured pruning. As the first work in implementing structured pruning on deep SNNs, we compare our method with the SOTA methods for ANN pruning NSlim (Liu et al., 2017) and GReg (Wang et al., 2021), reporting the pruned test accuracy and the float point operations reduction (FLOPs $\downarrow$ in Table 4) in inference. Additionally, we take NSlim as our baseline because our method develops based on it. Our method outperforms GReg (Wang et al., 2021), which is more complex and computationally expensive. It also shows significant improvements compared to the baseline, all while maintaining the equivalent level of FLOPs.
5.3 Effect of Neuron Criticality
We compare the baseline and pruning based on neuron criticality on CIFAR100. The results of the ablation study are shown in Table 5, where the proposed method outperforms the baseline at all sparsity levels. Furthermore, to evaluate the contribution of the criticality-based regeneration mechanism, we perform a performance analysis in Fig. 2 using various regeneration metrics. “Criticality” refers to the method based on criticality. “Grad” employs parameter gradients at the current iteration. “Random” adopts random ranking, and “GraNet” represents the GraNet gradient regeneration strategy (Liu et al., 2021). We observe that criticality-based regeneration consistently outperforms Grad and Random regeneration, highlighting our method’s superiority.
Figure 2: Comparison of regeneration methods based on criticality (ours), gradient, random, and GraNet (Liu et al., 2021). The two dashed lines represent the performance of the dense model and the baseline, respectively. (a) Accuracy of VGG16 with 90% sparsity for CIFAR100. (b) Accuracy of ResNet19 with 90% sparsity for CIFAR100.
5.4 Effect of Regeneration Ratio
To evaluate the robustness of the criticality-based regeneration, we conduct performance analysis in Fig. 2 using different metrics with different regeneration ratios $r$. The two dashed lines represent the dense model and baseline performance, respectively. We observe that the criticality-based method exhibits stable performance within the ratio range from 0.1 to 0.6 and higher accuracy than the baseline and other methods, indicative of the robustness over the wide range of $r$. GraNet exhibits different performance dynamics compared to other methods. We attribute this to the cosine decay of the regeneration ratio, gradually decreasing the impact of GMP as training progresses.
Sparsity 90.00% 95.69% 98.13%
@@ 5.4 Effect of Regeneration Ratio
5.5 Effect of Surrogate Function
To evaluate the effectiveness and robustness of the criticality mechanism for different surrogate functions, we report the pruning performance of different surrogate functions on VGG16 for CIFAR100. As shown in Fig. 3(a), we compare three surrogate functions, tanh, sigmoid, and arc tangent, for the dense model, baseline, and proposed method. The proposed method achieves higher performance compared to the baseline across different surrogate functions. The formulas of tanh, sigmoid, and arc tangent functions are respectively represented as Eq. 17, Eq. 18, Eq. 2a.
$g_{1}(x)=\frac{1}{2}\cdot\frac{tanh(3x)}{tanh(1.5)}+\frac{1}{2},$ (17)
$g_{2}(x)=\text{sigmoid}(4x)=\frac{1}{1+e^{-4x}},$ (18)
6 Discussion
In order to explore the underlying impact of neuron criticality during pruning and fine-tuning, we present the experimental results and statistical findings for the pruned model in this section for more in-depth analysis.
6.1 Importance Transition
We compare the importance transition in non-overlapping surviving structures between the models obtained by our method and the baseline. To better distinguish non-overlapping structures, we focus on the results of structured pruning. In Fig. 3(b), we present the means of $\gamma$ (normalized) of non-overlapping channels in the models through the fine-tuning process with L1 sparsity regularization for two methods. The results show that the channels regenerated by our method exhibit significantly higher importance than those without regeneration. Considering that the regenerated channels initially corresponded to lower $\gamma$ values and higher criticality, this result suggests the following role of our method: identifying structures with potential, enabling the high-criticality structures to occupy key positions and efficiently leading to higher performance gains.
@@ 6.1 Importance Transition
Figure 5: Comparison of the difference in feature extraction of models pruned by our method and the baseline. Our method achieves more uniform feature representations. (a) Intra-cluster variance of VGG-16 through unstructured pruning. (b) Intra-cluster variance of ResNet-19 through structured pruning. (c) Cosine similarity of the means of features for each class between the training and test dataset on VGG-16 through unstructured pruning. (d) Cosine similarity of the means of features for each class between the training and test dataset on ResNet-19 through structured pruning.
Figure 6: (a) Train and test loss (normalized) on VGG16 for CIFAR100 after unstructured pruning using the proposed method and the baseline. (b) Train and test loss (normalized) on ResNet19 for CIFAR100 after structured pruning using the proposed method and the baseline.
@@ 6.2 Feature Extraction
We compare the difference in feature map of models pruned by our method and the baseline. Fig. 4(a), 5(a) and 5(b) show the intra-cluster (I-C) variance (Kiang, 2001) of the features in CIFAR100 for unstructured and structured pruning. The intra-cluster variance measures the compactness of intra-cluster sample representations (Oti et al., 2020). And it computes the sum of squared distance between cluster samples and the corresponding centroid:
[…]
where $\mathbf{\mu}_{k}$ is the mean vector.
In our experiment, we divide intra-cluster variance by the number of samples in the cluster for the comparison. We extract feature maps before the fully connected classifier and normalize them to eliminate the influence of absolute values. According to Fig. 4(a), our model learns more consistent feature representations and achieves a lower mean of variances, indicating enhanced compactness in class features. Fig. 5(a) and 5(b) further indicate that our model achieved lower variances in 93% of classes on average. This improvement is consistent for both unstructured and structured pruning.
Furthermore, Fig. 4(b) illustrates the mean of cosine similarity of the class feature map between the training and test datasets. Fig. 5(c) and 5(d) further illustrate the cosine similarity of sample features in over 98% of classes. Our model exhibits higher similarity across almost all classes, leading to a more uniform representation of features between the train and test samples and reducing noise. Fig. 4(c) and 4(d) show the accuracies of pruned models for the training and test set after unstructured and structured pruning. Our model exhibits higher test accuracy after pruning but lower train accuracy than the baseline. We also report the training and test loss of pruned models by our method and the baseline in Fig. 6, and the results demonstrate a similar pattern to Fig. 4(c) and 4(d).
Based on the above observations, the baseline models suffer from overfitting during fine-tuning. The pruned models overfit the training set features, resulting in a drop in inference performance. Our method mitigates this issue by enabling the critical model to learn more consistent feature representations across the training and test sets. As a result, it reduces the noise between training and test samples, leading to improved inference performance.